While using a UV-Vis spectroscopy for concentration of the analyte of 0.11 and 0.21 mM, the absorbance read are 2 and 3 units receptively. What is the concentration if the absorbance read is 2.6 units
1
Expert's answer
2020-03-25T11:01:05-0400

Assuming linear dependence of concentration C on the absorbance A, we can write the linear equation:


"\frac{C-C_1}{C_2-C_1}=\frac{A-A_1}{A_2-A_1},\\
\space\\
\frac{C-0.11}{0.21-0.11}=\frac{A-2}{3-2},"

express C(A):


"C(A)=0.1A-0.09,\\
C(2.6)=0.17\text{ mM}."